# Decomposing Multivariate Information Rates in Networks of Random Processes

### v2

@@ II.2 Coarse-grained PID
$I(T;S)=\sum_{m=1}^{M}\mathcal{U}(T;S_{m})+\mathcal{R}(T;S)+\mathcal{S}(T;S),$ (7)
@@
$\mathcal{U}(T;S_{m})=I(T;S_{m})-\mathcal{R}(T;S).$ (8)
This approach provides a coarse-grained decomposition with a small number of atoms that scale gracefully with the system size: while the full PID (2) yields a number of atoms $|\mathcal{A}|$ that grows super-exponentially with $M$ like the Dedekind numbers Gutknecht et al. (2021), the coarse-grained PID (7) decomposes the multivariate information into exactly $M+2$ quantities. In particular, the two formulations coincide when $M=2$ source variables are considered, yielding $\mathcal{R}(T;S)=I^{\delta}(T;S_{\{1\}\{2\}})$, $\mathcal{U}(T;S_{1})=I^{\delta}(T;S_{\{1\}})$, $\mathcal{U}(T;S_{2})=I^{\delta}(T;S_{\{2\}})$, and $\mathcal{S}(T;S)=I^{\delta}(T;S_{\{12\}})$ (Fig. 1a, right). On the other hand, when $M\geq 3$ the coarse-graining is implemented by summing the PI of some of the atoms in (2). This issue has been addressed in Rosas et al. (2020), where the construction for $M=2$ was generalized to $M=3$ sources through the so-called $k^{\mathrm{th}}$ order coarse-grained PID, which preserves the intuitive meaning that synergy, redundancy, and unique information have for $M=2$ sources (Fig. 1b, right). Specifically, considering $k=1$, the $1^{\mathrm{st}}$-order synergy, $\mathcal{S}(T;S)$, corresponds to the information about the target that is provided by the whole $S$ but is not contained in any subset of sources when considered separately from the rest (atoms surrounded by the blue shade in Fig. 1). Similarly, the $1^{\mathrm{st}}$-order redundancy, $\mathcal{R}(T;S)$, is the information held by at least two different groups of size $1$ (magenta in Fig. 1). Finally, the $1^{\mathrm{st}}$-order unique information provided by the $m^{\mathrm{th}}$ source, $\mathcal{U}(T;S_{m})$, $m=1,\ldots,M$, is the information that $S_{m}$ has access to and no other subset of parts has access to on its own, although bigger groups of other parts may have (gray in Fig. 1).
@@ III Decomposition of multivariate information rates
Let us consider a dynamic network system composed of $M+1$ nodes, $\mathcal{Z}=\{\mathcal{Y},\mathcal{X}_{1},\ldots,\mathcal{X}_{M}\}$, where the activity at each node is described in terms of the vector random process $Z=\{Y,X_{1},\ldots,X_{M}\}$, with $Y$ assumed as target and $X=\{X_{1},\ldots,X_{M}\}$ as the vector of sources. Here we consider discrete-time random processes intended as time-ordered vector random variables: e.g., $Y_{n}\equiv Y(t_{n})$ is the variable sampling the target process at the time $t_{n}$, where $n\in\mathbb{N}$ is the time index; typically, $t_{n}=n\Delta t$, where $\Delta t$ is the sampling period ($f_{s}=\Delta t^{-1}$ is the sampling frequency). Since each random process is a collection of random variables, the application of the PID in the dynamic case is not straightforward as it implies an arbitrary selection of the variables to be extracted from each process. The most intuitive choice is to apply the PID to the variables sampling the processes at the same time $t_{n}$, setting $T=Y_{n}$ and $S=X_{i}$ as target and source variables and thus decomposing the static MI $I(Y_{n};X_{n})$ through a redundancy function based on zero-lag MI terms (e.g., $I^{\cap}(T;S_{i})=I(Y_{n};X_{i,n})$ according to the formalism presented in Sect. "Partial Information Decomposition and the redundancy lattice". An alternative approach is to decompose the joint transfer entropy (TE) from all sources to the target, which is defined as $T_{X\rightarrow Y}=I(Y_{n};X_{<n}|Y_{<n})$ Schreiber (2000), where $X_{<n}=\{X_{n-1},X_{n-2},\ldots\}$ and $Y_{<n}=\{Y_{n-1},Y_{n-2},\ldots\}$ denote the (potentially infinite-dimensional) vectors representing the past history of the source and target processes; in this case the PID is applied setting $T=Y_{n}$ and $S=X_{<n}$ as target and source variables, and the redundancy function is a conditional MI (e.g., $I^{\cap}(T;S_{i})=I(Y_{n};X_{i,<n}|Y_{<n}))$. Although these two alternative applications of PID to random processes are common Kay et al. (2022); Varley et al. (2023a); Koçillari et al. (2023); Luppi et al. (2024a), they provide only a partial view of the dynamic interactions among the processes and are based on implicit assumptions which are often not fulfilled by dynamic network systems. The zero-lag PID of $I(Y_{n};X_{n})$ presupposes to work with memoryless processes, a condition that is typically not satisfied in practice as the presence of a temporal statistical structure is inherently expected in time series data. On the other hand, the PID of the TE $I(Y_{n};X_{<n}|Y_{<n})$ cannot account for instantaneous interactions among the processes, nor for causal interactions occurring in the causal direction from the target to the sources. To overcome these limitations, in the following, we propose a framework for the decomposition of the information shared dynamically between the target and the source processes.
@@ III.1 Partial Information Rate Decomposition
The framework proposed here for the decomposition of information in multivariate random processes makes use of the concepts of entropy rate (ER) and mutual information rate (MIR). The ER of a generic stationary random processes $X$ is defined as Cover (1999)
$H_{X}=\lim_{m\to\infty}\frac{1}{m}H(X_{n:n+m}),$ (11)
@@
and, being equivalent to the conditional entropy of the present state of the process given its past states (i.e., $H_{X}=H(X_{n}|X_{<n})$), quantifies the rate of generation of new information in the process. Then, considering another process $Y$, the MIR between $X$ and $Y$ is Duncan (1970)
$I_{X;Y}=\lim_{m\to\infty}\frac{1}{m}I(X_{n},\ldots,X_{n+m};Y_{n},\ldots,Y_{n+m}),$ (12)
@@
quantifying the information shared by the two processes per unit of time; the MIR can be expressed in terms of entropy rates as $I_{X;Y}=H_{X}+H_{Y}-H_{X,Y}$, evidencing the analogy between the concepts of entropy and MI for random variables and the concepts of ER and MIR for random processes.
@@
Exploiting this analogy, we use the MIR as a building block for assessing and decomposing the dynamic information shared between the scalar target process $Y$ and the vector source process $X=\{X_{1},\ldots,X_{M}\}$ of the analyzed network system. Specifically, we formalize a so-called Partial Information Rate Decomposition (PIRD) which makes use of the same lattice structure of the PID Williams and Beer (2010) to expand the MIR between target and sources as:
$I_{Y;X}=\sum_{\alpha\in\mathcal{A}}I^{\delta}_{Y;X_{\alpha}},$ (13)
@@
where $I^{\delta}_{\cdot;\cdot}$ is a PI rate function defined for each atom $\alpha=\{\alpha_{1},\ldots,\alpha_{J}\}$ of the lattice, and $X_{\alpha}$ denotes the $\alpha^{\mathrm{th}}$ set of subsets of source processes, with $X_{\alpha_{j}}\subseteq X$. As happens for the PID, to solve the PIRD it is necessary to define a so-called redundancy rate function, here denoted as $I^{\cap}_{\cdot;\cdot}$, which generalizes the MIR over the lattice and replaces the concept of redundancy function generalizing the MI. The redundancy rate of the $\alpha^{\mathrm{th}}$ atom is obtained summing the PI rate of the same atom to the PI rates of the atoms positioned downwards in the lattice:
$I^{\cap}_{Y;X_{\alpha}}=\sum_{\beta\preceq\alpha}I^{\delta}_{Y;X_{\beta}};$ (14)
@@
then, once the redundancy rate is known, the information rate associated to all atoms can be retrieved via Möbius inversion of (14). In Sect. "Frequency-domain PIRD" we will elaborate on the definition of redundancy rate functions.
@@
Importantly, since the PIRD is built over the same lattice backbone as the PID, several concepts and relations defined for the PID still hold in the dynamic case (see Fig. 1, with the shrewdness of considering information rates in place of the static information distributed over the lattice). For instance, as the PIRD satisfies the same consistency equations valid for the PID, the redundancy function computed for an atom composed by one single source reduces to the MIR between the target and that source, i.e. $I^{\cap}_{Y;X_{\alpha}}=I_{Y;X_{i}}$ when $\alpha=\{i\}$. Moreover, the rate of information shared between the $M$ source processes $X_{1},\ldots,X_{M}$ taken together and the target process $Y$ can be expanded in analogy to (7) as the sum of $M+2$ contributions:
$I_{Y;X}=\sum_{m=1}^{M}\mathcal{U}_{Y;X_{m}}+\mathcal{R}_{Y;X}+\mathcal{S}_{Y;X};$ (15)
@@
eq. (15) achieves a so-called coarse-grained PIRD, whereby each of the $M$ terms $\mathcal{U}_{Y;X_{m}}=I_{Y;X_{m}}-\mathcal{R}_{Y;X}$ identifies the unique rate of information produced by $Y$ that is shared exclusively with $X_{m}$ ($m=1,\ldots,M$), the term $\mathcal{R}_{Y;X}$ identifies the redundant rate of information produced by $Y$ that is shared simultanously with all the source processes in $X$, and the term $\mathcal{S}_{Y;X}$ identifies the synergistic rate of information produced by $Y$ that only arises from knowing all the sources $X_{1},\ldots,X_{M}$. These coarse-grained information rates correspond for the case of two sources to the PI rate of the four atoms of the redundancy rate lattice (Fig. 1a, right), while they can be obtained for the case $M=3$ by summing the PI rates of the atoms with redundant, unique or synergistic character (respectively, magenta, gray and blue shades in Fig. 1b, right).
@@
As a further remark, we stress that the PIRD formulated in (13) decomposes the information rates shared by multivariate processes accounting for their full dynamical structure and, as such, it generalizes previous attempts to apply the PID to random processes Kay et al. (2022); Varley et al. (2023a); Koçillari et al. (2023); Luppi et al. (2024a). This aspect becomes apparent considering that the MIR (12) can be expanded as the sum of three terms, two related to the causal (time-lagged) information transfer along the two directions $X\rightarrow Y$ and $Y\rightarrow X$, and the latter related to the instantaneous (zero-lag) information shared between $X$ and $Y$ (see, e.g., Barà et al. (2023)):
$I_{X;Y}=T_{X\rightarrow Y}+T_{Y\rightarrow X}+I_{X\cdot Y},$ (16)
@@
where $T_{X\rightarrow Y}=I(Y_{n};X_{<n}|Y_{<n})$ and $T_{Y\rightarrow X}=I(X_{n};Y_{<n}|X_{<n})$ are the transfer entropies measuring causal information transfer, and $I_{X\cdot Y}=I(X_{n};Y_{n}|X_{<n},Y_{<n})$ is the instantaneous information shared between the processes. Now, if we consider the case in which the overall multivariate process $Z=\{X,Y\}$ is memoryless, meaning that its whole temporal statistical structure is absent, the two TEs vanish and the instantaneous term becomes the information shared instantaneously between $X$ and $Y$, showing how the PIRD reduces to a PID applied to the static MI $I(X_{n};Y_{n})$. If, on the other hand, we consider the case of strictly causal processes with exclusive (unidirectional) information transfer from the sources to the target, both the instantaneous term and the TE from $Y$ to $X$ vanish, and the PIRD reduces to a PID applied to the TE $T_{X\rightarrow Y}$. These properties will be illustrated in simulated examples in Sect. "Theoretical Examples".

### v3

@@ II.2 Coarse-grained PID
$I(T;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{S}})=\sum_{m=1}^{M}\mathcal{U}(T;S_{m})+\mathcal{R}(T;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{S}})+\mathcal{S}(T;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{S}}),$ (7)
@@
$\mathcal{U}(T;S_{m})=I(T;S_{m})-\mathcal{R}(T;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{S}}).$ (8)
This approach provides a coarse-grained decomposition with a small number of atoms that scale gracefully with the system size: while the full PID (2) yields a number of atoms $|\mathcal{A}|$ that grows super-exponentially with $M$ like the Dedekind numbers Gutknecht et al. (2021), the coarse-grained PID (7) decomposes the multivariate information into exactly $M+2$ quantities. In particular, the two formulations coincide when $M=2$ source variables are considered, yielding $\mathcal{R}(T;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{S}})=I^{\delta}(T;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{S}}_{\{1\}\{2\}})$, $\mathcal{U}(T;S_{1})=I^{\delta}(T;S_{\{1\}})$, $\mathcal{U}(T;S_{2})=I^{\delta}(T;S_{\{2\}})$, and $\mathcal{S}(T;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{S}})=I^{\delta}(T;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{S}}_{\{12\}})$ (Fig. 1a, right). On the other hand, when $M\geq 3$ the coarse-graining is implemented by summing the PI of some of the atoms in (2). This issue has been addressed in Rosas et al. (2020), where the construction for $M=2$ was generalized to $M=3$ sources through the so-called $k^{\mathrm{th}}$ order coarse-grained PID, which preserves the intuitive meaning that synergy, redundancy, and unique information have for $M=2$ sources (Fig. 1b, right). Specifically, considering $k=1$, the $1^{\mathrm{st}}$-order synergy, $(T;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{S}})$ S ), corresponds to the information about the target that is provided by the whole S but is not contained in any subset of sources when considered separately from the rest (atoms surrounded by the blue shade in Fig. 1). Similarly, the $1^{\mathrm{st}}$-order redundancy, $\mathcal{R}(T;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{S}})$, is the information held by at least two different groups of size $1$ (magenta in Fig. 1). Finally, the $1^{\mathrm{st}}$-order unique information provided by the $m^{\mathrm{th}}$ source, $\mathcal{U}(T;S_{m})$, $m=1,\ldots,M$, is the information that $S_{m}$ has access to and no other subset of parts has access to on its own, although bigger groups of other parts may have (grey in Fig. 1).
@@ III Decomposition of multivariate information rates
Let us consider a dynamic network system composed of $M+1$ nodes, $\mathcal{Z}=\{\mathcal{Y},\mathcal{X}_{1},\ldots,\mathcal{X}_{M}\}$, where the activity at each node is described in terms of the vector random process ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{Z}}=\{Y,X_{1},\ldots,X_{M}\}$, with $Y$ assumed as target and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}=\{X_{1},\ldots,X_{M}\}$ as the vector of sources. Here we consider discrete-time random processes intended as time-ordered vector random variables: e.g., $Y(t_{n})$ is the variable sampling the target process at the time $t_{n}$, where $n\in\mathbb{Z}$ is the time index; typically, $t_{n}=n\Delta t$, where $\Delta t$ is the sampling period ($f_{s}=\Delta t^{-1}$ is the sampling frequency); see Fig. 2 for a graphical representation.
Figure 2: Selection of the variables to be used in PID schemes applied to a network of random processes. Each panel depicts the random variables (circles) sampling the vector random process $\{Y,\textbf{X}\}=\{Y,X_{1},\ldots,X_{M}\}$ at specific time points. a) The zero-lag PID is applied to the MI between variables sampling the processes at the same time $t_{n}$, taking $Y(t_{n})$ (blue) and $\textbf{X}(t_{n})$ (magenta) as target and source variables. b) The PID applied to the joint TE considers the variable sampling the process $Y$ at the current time $t_{n}$, $Y(t_{n})$, as target variable (blue) and the variables sampling the past of the process X at times preceding $t_{n}$, $\textbf{X}(t_{<n})$, as the set of sources (magenta), and decomposes a conditional MI where the past of $Y$, $Y(t_{<n})$, is used for conditioning (pink). c) The PIRD considers the whole processes as building blocks for the decomposition, and is applied to the MIR between the target process $Y$ (blue) and the source processes in X (magenta).
@@
Since each random process is a collection of random variables, the application of PID in the dynamic case is not straightforward as it implies an arbitrary selection of the variables to be extracted from each process. The most intuitive choice is to apply the PID to the variables sampling the processes at the same time $t_{n}$, setting $T=Y(t_{n})$ and $\textbf{S}=\textbf{X}(t_{n})$ as target and source variables and thus decomposing the static MI $I(Y(t_{n});\textbf{X}(t_{n}))$ through a redundancy function based on zero-lag MI terms (e.g., $I^{\cap}(T;S_{i})={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}I(Y(t_{n});X_{i}(t_{n}))}$) according to the formalism presented in Sect. "Partial Information Decomposition and the redundancy lattice" (Fig. 2a). An alternative approach is to decompose the joint transfer entropy (TE) from all sources to the target, which is defined as $T_{\textbf{X}\rightarrow Y}=I(Y(t_{n});\textbf{X}(t_{<n})|Y(t_{<n}))$ Schreiber (2000), where $\textbf{X}(t_{<n})=\{\textbf{X}(t_{n-1}),\textbf{X}(t_{n-2}),\ldots\}$ and $Y(t_{<n})=\{Y(t_{n-1}),Y(t_{n-2}),\ldots\}$ denote the (potentially infinite-dimensional) vectors representing the past history of the source and target processes; in this case the PID is applied setting $T=Y(t_{n})$ and $\textbf{S}=\textbf{X}(t_{<n})$ as target and source variables, and the redundancy function is a conditional MI (e.g., $I^{\cap}(T;S_{i})={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}I(Y(t_{n});X_{i}(t_{<n})|Y(t_{<n}))}$, Fig. 2b).
Although these two alternative applications of PID to random processes are common Kay et al. (2022); Varley et al. (2023a); Koçillari et al. (2023); Luppi et al. (2024a), they provide only a partial view of the dynamic interactions among the processes and are based on implicit assumptions which are often not fulfilled by dynamic network systems. The zero-lag PID of $I(Y(t_{n});\textbf{X}(t_{n}))$ presupposes to work with memoryless processes, a condition that is typically not satisfied in practice as the presence of a temporal statistical structure is inherently expected in time series data. On the other hand, the PID of the TE $I(Y(t_{n});\textbf{X}(t_{<n})|Y(t_{<n}))$ cannot account for instantaneous interactions among the processes, nor for causal interactions occurring in the causal direction from the target to the sources. To overcome these limitations, in the following we propose a framework for the decomposition of the information shared dynamically between the target and the source processes.
@@ III.1 Partial Information Rate Decomposition
The framework proposed here for the decomposition of information in multivariate random processes makes use of the concepts of entropy rate (ER) and mutual information rate (MIR). The ER of a generic stationary vector random processes X is defined as Cover (1999)
$H_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}}=\lim_{m\to\infty}\frac{1}{m}H\big{(}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}(t_{n:n+m})}\big{)},$ (11)
@@
where $\textbf{X}(t_{n:n+m})=\{\textbf{X}(t_{n}),\ldots,\textbf{X}(t_{n+m})\}$. Being equivalent to the conditional entropy of the present state of the process given its past states (i.e., $H_{\textbf{X}}=H(\textbf{X}(t_{n})|\textbf{X}(t_{<n}))$), the ER quantifies the rate of generation of new information in the process. Then, considering another process $Y$, the MIR between X and $Y$ is defined as Duncan (1970)
$I_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}};Y}=\lim_{m\to\infty}\frac{1}{m}I\big{(}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}(t_{n:n+m});Y(t_{n:n+m})}\big{)},$ (12)
@@
quantifying the information shared by the two processes per unit of time; the MIR can be expressed in terms of entropy rates as $I_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}};Y}=H_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}}+H_{Y}-H_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}},Y}$, evidencing the analogy between the concepts of entropy and MI for random variables and the concepts of ER and MIR for random processes.
@@
Exploiting this analogy, we use the MIR as a building block for assessing and decomposing the dynamic information shared between the scalar target process $Y$ and the vector source process ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}=\{X_{1},\ldots,X_{M}\}$ of the analyzed network system (Fig. 2c). Specifically, we formalize a so-called Partial Information Rate Decomposition (PIRD) which makes use of the same lattice structure of the PID Williams and Beer (2010) to expand the MIR between target and sources as:
$I_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}}=\sum_{\alpha\in\mathcal{A}}I^{\delta}_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\alpha}},$ (13)
@@
where $I^{\delta}_{\cdot;\cdot}$ is a PI rate function defined for each atom $\alpha=\{\alpha_{1},\ldots,\alpha_{J}\}$ of the lattice, and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\alpha}$ denotes the $\alpha^{\mathrm{th}}$ set of subsets of source processes, with $_{\alpha_{j}}\subseteq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}$ X. As happens for the PID, to solve the PIRD it is necessary to define a so-called redundancy rate function, here denoted as $I^{\cap}_{\cdot;\cdot}$, which generalizes the MIR over the lattice and replaces the concept of redundancy function generalizing the MI. The redundancy rate of the $\alpha^{\mathrm{th}}$ atom is obtained summing the PI rate of the same atom to the PI rates of the atoms positioned downwards in the lattice:
$I^{\cap}_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\alpha}}=\sum_{\beta\preceq\alpha}I^{\delta}_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\beta}};$ (14)
@@
then, once the redundancy rate is known, the information rate associated to all atoms can be retrieved via Möbius inversion of (14). In the Section "Frequency-domain PIRD" we will elaborate on the definition of redundancy rate functions.
@@
Importantly, since the PIRD is built over the same lattice backbone as the PID, several concepts and relations defined for the PID still hold in the dynamic case (see Fig. 1, with the shrewdness of considering information rates in place of the static information distributed over the lattice). For instance, as the PIRD satisfies the same consistency equations valid for the PID, the redundancy function computed for an atom composed by one single source reduces to the MIR between the target and that source, i.e. $I^{\cap}_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\alpha}}=I_{Y;X_{i}}$ when $\alpha=\{i\}$. Moreover, the rate of information shared between the $M$ source processes $X_{1},\ldots,X_{M}$ taken together and the target process $Y$ can be expanded in analogy to (7) as the sum of $M+2$ contributions:
$I_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}}=\sum_{m=1}^{M}\mathcal{U}_{Y;X_{m}}+\mathcal{R}_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}}+\mathcal{S}_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}},$ (15)
@@
achieving a so-called coarse-grained PIRD whereby each of the $M$ terms $\mathcal{U}_{Y;X_{m}}=I_{Y;X_{m}}-\mathcal{R}_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}}$ identifies the unique rate of information produced dynamically by $Y$ that is shared exclusively with $X_{m}$ ($m=1,\ldots,M$), the term $\mathcal{R}_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}}$ identifies the redundant rate of information that is shared simultaneously with all the source processes in X, and the term $\mathcal{S}_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}}$ identifies the synergistic rate of information that only arises from knowing all the source processes $X_{1},\ldots,X_{M}$. Moreover, a quantity reflecting the dynamic redundancy-synergy balance can be obtained as $\Delta_{Y;\textbf{X}}=\mathcal{R}_{Y;\textbf{X}}-\mathcal{S}_{Y;\textbf{X}}$. These coarse-grained information rates correspond for the case of two sources to the PI rate of the four atoms of the redundancy rate lattice (Fig. 1a, right), while they can be obtained for the case $M=3$ by summing the PI rates of the atoms with redundant, unique or synergistic character (respectively, magenta, grey and blue shades in Fig. 1b, right).
@@
As a further remark, we stress that the PIRD formulated in (13) decomposes the information rates shared by multivariate processes accounting for their full dynamical structure and, as such, it generalizes previous attempts to apply the PID to random processes Kay et al. (2022); Varley et al. (2023a); Koçillari et al. (2023); Luppi et al. (2024a). This aspect becomes apparent considering that the MIR (12) can be expanded as the sum of three terms, two related to the causal (time-lagged) information transfer along the two directions ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}\rightarrow Y$ and $Y\rightarrow{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}$, and the latter related to the instantaneous (zero-lag) information shared between X and $Y$ (see, e.g., Barà et al. (2023)):
$I_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}};Y}=T_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}\rightarrow Y}+T_{Y\rightarrow{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}}+I_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}\cdot Y},$ (16)
@@
where $T_{\textbf{X}\rightarrow Y}=I(Y(t_{n});\textbf{X}(t_{<n})|Y(t_{<n}))$ and $T_{Y\rightarrow\textbf{X}}=I(\textbf{X}(t_{n});Y(t_{<n})|\textbf{X}(t_{<n}))$ are the transfer entropies measuring causal information transfer, and $I_{\textbf{X}\cdot Y}=I(\textbf{X}(t_{n});Y(t_{n})|\textbf{X}(t_{<n}),Y(t_{<n}))$ is the instantaneous information shared between the processes. Now, if we consider the case in which the overall multivariate process ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{Z}}=\{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}},Y\}$ is memoryless, meaning that its whole temporal statistical structure is absent, the two TEs vanish and the instantaneous term becomes the information shared instantaneously between X and $Y$, showing how the PIRD reduces to a PID applied to the static MI $I(\textbf{X}(t_{n});Y(t_{n}))$. If, on the other hand, we consider the case of strictly causal processes with exclusive (unidirectional) information transfer from the sources to the target, both the instantaneous term and the TE from $Y$ to X vanish, and the PIRD reduces to a PID applied to the TE $T_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}\rightarrow Y}$. These properties will be illustrated in simulated examples in Sect. "Theoretical Examples".